If 0.100 mole of AgNO3(s) is added to 300 mL of a 2.00 M NH3(aq) solution, what is [Ag+] in the final solution?
1
Expert's answer
2020-04-23T13:08:31-0400

"c=n\/V" , c - solution molarity, n - chemical amount of substance, V - solution volume (L).

The concentration of AgNO3 is 0,100mol/0,3L=0,333mol/L.The concentration of Ag+=0,333mol/L.

Silver forms a complex with ammonia: "Ag^+ +2NH_3 \to [Ag(NH_3)_2 ]^+"

Stability constant of this complex is "K=[Ag(NH_3)_2]^+\/[Ag]*[NH_3]^2 =" 1,075*107

The dissociation of ammonia is neglected. Let the concentration of the complex is x, then the concentration of Ag+ is 0,333-x. The concentration of ammonia should be (2,00-x)2 but the the resulting equation is difficult, so we can neglect this momemt and resulting equation is

"1,075*10^7 =x\/(2,00^2*(0,333-x))"

express x:

"x=1,43*10^7-4,3*10^7x"

"x=0,3326" mol/L

[Ag+]=0,333-0,3326=4*10-4mol/L

Answer:4*10-4mol/L